OpenStudy (anonymous):

Which of the following sets could be the sides of a right triangle? {2, 3, square root of 10} {3, 5, square root of 34} {5, 8, 12}

OpenStudy (anonymous):

and why?

OpenStudy (anonymous):

a^2 + b^2 = c^2 So {3, 5, root 34} since 3^2 + 5^2 = 34